The following table shows the income distribution of 600 families. Find the minimum income


of the riches 30% families. Also the limits of income of middle 50% of families, to the nearest


rupees.


Income Below


75


75-


150


150-


225


225-


300


300-


375


375-


400


400 &


above


No. of


families


69 137 225 46 88 25 10


Ans.: the richest 30 % families earns Rs. 222 and above per week , the middle 50% families


weekly income lies between 120 and 256.


1
Expert's answer
2022-01-25T16:47:48-0500

Class f cf

0-75 69 69

75-150 137 206

150-225 225 431

225-300 46 477

300-375 88 565

375-400 25 590

400+ 10 600

To find the minimum income of the riches 30% families, we determine the "70^{th}" percentile given as,

"P_{70}=l+({70\times n\over 100}-cf)\times{c\over f}" where, "n=600" and

"l" is the lower class boundary of the class containing "P_{70}"

"f" is the frequency of the class containing "P_{70}"

"c" is the width of the class containing "P_{70}"

"cf" is the cumulative frequency of the class preceding the class with "P_{70}".

The class containing "P_{70}" is,

"({70\times n\over 100})^{th} class=({70\times 600\over 100})=420". Therefore the class with "P_{70}" is 150-225

Thus,

"P_{70}=150+(420-206)\times {75\over 225}=150+71.33=221.33\approx 222"


The limits of income of middle 50% of families is same as determining "Q_1" and "Q_3" where,

"Q_1" is the lower limit and "Q_3" is the upper limit.

Therefore,

"Q_1=l+({n\over4}-cf)\times {c\over f}" where,

"l" is the lower class boundary of the class containing "Q_1"

"f" is the frequency of the class containing "Q_1"

"c" is the width of the class containing "Q_1"

"cf" is the cumulative frequency of the class preceding the class with "Q_1".

The class containing "Q_1" is,

"({n\over 4})^{th} class={600\over 4}=150".Therefore, the class with "Q_1" is 75-150

Thus,

"Q_1=75+(150-69)\times {75\over 137}=119.34\approx 120"


and,

"Q_3=l+({3n\over4}-cf)\times {c\over f}" where,

"l" is the lower class boundary of the class containing "Q_3"

"f" is the frequency of the class containing "Q_3"

"c" is the width of the class containing "Q_3"

"cf" is the cumulative frequency of the class preceding the class with "Q_3".

The class containing "Q_1" is,

"({3n\over 4})^{th} class={1800\over 4}=450".Therefore, the class with "Q_3" is 150-225

Thus,

"Q_3=225+(450-431)\times {75\over 46}=255.98\approx 256"

Therefore, the limits of income of middle 50% of families is 120 and 256.